Amoxicillin 20 mg/kg/day divided q 8hr. is ordered for an 11 year old child with an ear, nose and throat infection. The child weighs 36 kg. Your bottle of amoxicillin is labeled 200 mg / 5 mL. How many mL of amoxicillin will you give the child at each dose? .
Answer #1

ANSWER.

* Calculation of ml of drug required in each dose.

       Formula

           Volume(ml) = Desired dose/Dose in hand * Quantity.

   Here

         Ordered dose = 20 mg/kg/day.

        Weight of the patient = 36 kg

      Dose required for the patient = 20 * 36

                                                     = 720 mg.

So

       Desired dose(D) = 720 mg/day.

       Dose in hand (H) = 200 mg

            Quantity (Q) = 5 ml

As per above formula

      Volume(ml) of drug required = D/H * Q

                                                    = 720/200 * 5

                                                    = 3.6 * 5

                                                    = 18 ml per day

      Volume(ml) of drug required per each dose = 18 ml divided q8hr or 3 times a day.

                                                                  = 18 / 3

                                                                  = 6 ml per dose.